How many ounces of a
38% alcohol solution and a
43% alcohol solution must be combined to obtain
65 ounces of a
42% solution?

Answer by stanbon(75887)

How many ounces of a
38% alcohol solution and a
43% alcohol solution must be combined to obtain
65 ounces of a
42% solution?
---------
Equation::
alc + alc = alc
0.38x + 0.43(65-x) = 0.42*65
--------
38x + 43*65 - 43x = 42*65
------
-5x = -1*65
-----
x = 13 oz (amt. of 38% solution needed)
65-x = 52 oz (amt. of 43% solution needed)
-------------
Cheers,
Stan H.
--------------